OpenStudy (anonymous):

(over:∞)Σ(under:n=1) (1/3)^n What makes this into (1/2)<∞ ? I dont get the connection.... Can anyone explain?

OpenStudy (anonymous):

is it \[\sum_{\infty}^{n=1} (\frac{ 1 }{ 3 })^{n}\] ...????

OpenStudy (anonymous):

other way around n=1 is under, but the rest, yes.

OpenStudy (anonymous):

Found it! :D \[\sum_{n=1}^{\infty} (\frac{ 1 }{ 3 })^{n} equals \sum_{n=0}^{\infty} (\frac{ 1 }{ 3 })*(\frac{ 1 }{ 3 })^{n} \] And that equals \[\frac{ \frac{ 1 }{ 3 } }{ 1-\frac{ 1 }{ 3 } } equals \frac{ 1 }{ 2 }\]